OpenStudy (anonymous):

the ratio of two numbers is 5 to 3. the sum of the numbers is 16. what are the numbers A8 AND 8 B9 AND 7 C15 AND 9 D 10 AND 6

OpenStudy (arbershabani97):

10 and 6 , 10/6=5/3 10+6=16 you just have to check which ones work

OpenStudy (anonymous):

Make two equations k(5) = n(3) and k+n = 16 solve second for k and insert in first

OpenStudy (anonymous):

or do what arbershabani97 did